A BOX CONTAINS 12 RED T-SHIRTS AND 18 BLACK T-SHIRTS.
FIND THE PROBABILITY OF RANDOMLY SELECTING 3 DIFFERENT BLACK T-SHIRTS IN SUCCESION.....

Answer by stanbon(75887)

A BOX CONTAINS 12 RED T-SHIRTS AND 18 BLACK T-SHIRTS.
FIND THE PROBABILITY OF RANDOMLY SELECTING 3 DIFFERENT BLACK T-SHIRTS IN SUCCESSION.....
============
Ans = (18/30)(17/29)(16/28) = 18C3/30C3 = 0.200985
==============================================
Cheers,
Stan H.